# Stepwise Decomposition and Dual-stream Focus: A Novel Approach for Training-free Camouflaged Object Segmentation

### v2

@@ 3.3.1. Region-constrained Dual-stream Visual Prompting
(10) $p_{i}^{f}$ =$=\{(x,y)\mid\mathcal{N}(H_{i}^{f}\mid B_{i}^{{}^{\prime}}~{}\text{or}~{}B_{i})[x,y]\geq 0.9\},$
(11) $p_{i}^{b}$ =$=\{(x,y)\mid\mathcal{N}(H_{i}^{b}\mid B_{i}^{{}^{\prime}}~{}\text{or}~{}B_{i})[x,y]\geq 0.9\},$
@@ 3.4. Self-Consistency Mask Selection
Existing methods (Hu et al., 2024a, b) often rely on iterative, cycle-generation strategies to produce segmentation masks, repeatedly integrating previous iteration outputs (e.g., heatmaps (Hu et al., 2024a) or masks (Hu et al., 2024b)) back into the original image. However, this iterative dependency typically incurs substantial computational overhead. Inspired by the concept of self-consistency (Wang et al., 2023; Knappe et al., 2024; Chen et al., 2024b) employed in large language models — where the consistency among multiple answers to identical or semantically similar questions is evaluated to enhance reliability — we propose the Self-Consistency Mask Selection mechanism to obtain robust segmentation predictions without iterative dependency efficiently.
@@
After obtaining multiple candidate masks, we select the most representative segmentation mask by evaluating their mutual consistency. Formally, the final selected mask index $i$ is determined by minimizing the difference between each individual mask $M_{i}$ and the mean mask computed across all repetitions:
(13) $i=\arg\min_{i}\left(\left|M_{i}-\frac{\sum_{i}(M_{1},\dots,M_{I})}{i_{\text{result}}}\right|\right),$
where $M_{i}$ denotes the mask generated by the $i$-th repetition. This consensus-based selection criterion ensures the final segmentation result exhibits enhanced robustness and accuracy, effectively mitigating uncertainty induced by stochastic text prompt generation.
@@ 4.3. Ablation Study
Effectiveness of DRVP Module
Table 3. Ablation experiment of the two main strategies of the DRVP module.
@@
Table 4. Ablation study on the influence of repetition number $I$ in the proposed Self-Consistency Mask Selection.
@@ 5. Conclusion
In this work, we introduce RDVP-MSD, a novel training-free test-time adaptation framework that explicitly resolves the semantic ambiguity arising from instance-specific text prompts generation and mitigates the semantic discrepancy as well as spatial separation encountered during instance-specific visual prompts extraction within task-generic promptable segmentation scenarios for camouflaged objects. Leveraging the proposed MSD-CoT, RDVP-MSD progressively refines instance-specific text prompts, while our RDVP independently obtains the instance-specific visual prompts within spatial constraints. Extensive experiments demonstrate that RDVP-MSD achieves state-of-the-art segmentation accuracy across COS standard benchmarks with substantially improved efficiency without any training or supervision, thus establishing a new paradigm for efficient and precise camouflaged object segmentation.

### v3

@@ 3.3.1. Region-constrained Dual-stream Visual Prompting
(6) $p_{i}^{f}$ $=\{(x,y)\mid\mathcal{N}(H_{i}^{f}\mid B_{i}^{{}^{\prime}}~\text{or}~B_{i})[x,y]\geq 0.9\},$
(7) $p_{i}^{b}$ $=\{(x,y)\mid\mathcal{N}(H_{i}^{b}\mid B_{i}^{{}^{\prime}}~\text{or}~B_{i})[x,y]\geq 0.9\},$
@@ 3.4. Self-Consistency Mask Selection
Existing methods (Hu et al., 2024a, b) often rely on iterative, cycle-generation strategies to produce segmentation masks, repeatedly integrating previous iteration outputs (e.g., heatmaps (Hu et al., 2024a) or masks (Hu et al., 2024b)) back into the original image. However, this iterative dependency typically incurs substantial computational overhead. Inspired by the concept of self-consistency (Wang et al., 2023; Chen et al., 2024b) employed in large language models — where the consistency among multiple answers to identical or semantically similar questions is evaluated to enhance reliability — we propose the Self-Consistency Mask Selection mechanism to obtain robust segmentation predictions without iterative dependency efficiently.
@@
After obtaining multiple candidate masks, we select the most representative segmentation mask by evaluating their mutual consistency. Formally, the final selected mask index $i^{\star}$ is determined by minimizing the difference between each individual mask $M_{i}$ and the mean mask computed across all repetitions:
(9) $i^{\star}=\arg\min_{i}\left(\left|M_{i}-\frac{\sum_{i}(M_{1},\dots,M_{I})}{I}\right|\right),$
where $M_{i}$ denotes the mask generated by the $i$-th repetition. This consensus-based selection criterion ensures the final segmentation result $M_{i^{\star}}$ exhibits enhanced robustness and accuracy, effectively mitigating uncertainty induced by stochastic text prompt generation.
@@ 4.3. Ablation Study
Effectiveness of RDVP Module
Table 3. Ablation experiment of the two main strategies of the RDVP module.
@@
Table 4. Ablation study on the influence of repetition number II in the proposed Self-Consistency Mask Selection.
@@ 5. Conclusion
In this work, we introduce RDVP-MSD, a novel training-free test-time adaptation framework that explicitly mitigates the semantic ambiguity arising from instance-specific text prompts generation and mitigates the semantic discrepancy as well as spatial separation encountered during instance-specific visual prompts extraction within task-generic promptable segmentation scenarios for camouflaged objects. Leveraging the proposed MSD-CoT, RDVP-MSD progressively refines instance-specific text prompts, while our RDVP independently obtains the instance-specific visual prompts within spatial constraints. Extensive experiments demonstrate that RDVP-MSD achieves state-of-the-art segmentation accuracy across COS standard benchmarks with substantially improved efficiency without any training or supervision, thus establishing a new paradigm for efficient and precise camouflaged object segmentation.
Acknowledgements.
This work is supported in part by the Science and Technology Innovation Plan of Shanghai Science and Technology Commission under grant No.22511106005. We appreciate the High Performance Computing Center of Shanghai University, and Shanghai Engineering Research Center of Intelligent Computing System for providing computing resources and technical support.